THE ALGEBRAIC COMBINATORICS RENCONTRE jointly with the 34th SÉMINAIRE LOTHARINGIEN DE COMBINATOIRE

Domaine Saint-Jacques, Saint-Nabor, F-67530 Ottrott, France

March 12, 1995 to Thursday, March 16, 1995

LIST OF PARTICIPANTS

Aberystwyth (1): Alun Morris. Amsterdam (1): Marc A. A. van Leeuwen. Bayreuth (4): Adalbert Kerber, Thomas Scharf, Axel Kohnert, Alfred Wassermann. Berlin (3): Günter M. Ziegler, Jürgen Richter-Gebert, Jörg Rambau. Bielefeld (8): Andreas Dress, Vincent Moulton, Joern Bornhoeft, Werner Terhalle, Soeren Perrey, Daniel Huson, Olaf Delgado, Oliver Goodman. Bordeaux (4): Mireille Bousquet-Melou, Alain Denise, Walter Wenzel, Sacha Zvonkin. Cagliari (2): Luigi Cerlienco, Francesco Piras. Copenhagen (1): Joern Olsson. Erlangen (3): Volker Strehl, Roman Koenig, Ulrike Sattler. Magdeburg (1): Christine Bessenrodt. Marne-la-Vallée/Paris/Rouen (9): Bernard Leclerc, Jacques Désarménien, Gérard Duchamp, Pierre-André Picon, Jean-Yves Thibon, Alain Lascoux, Daniel Krob, Sébastien Veigneau, Jean-Christophe Novelli. Strasbourg (6): Dominique Foata, Jiang Zeng, Yaacob Akiba Slama, Jean-Pierre Jouanolou, Guoniu Han, Raphaelle Supper. Stockholm (5): Anders Björner, Kimmo Eriksson, Henrik Eriksson, Johan Karlander, Svante Linusson. Wien (2): Christian Krattenthaler, Michael Schlosser.

Other participants (15): Anna Helversen-Pasotto, Luise Unger (Paderborn), Henri Gaudier (Valenciennes), Daniel Barsky (Paris), Heinz Lüneburg (Kaiserslautern), Petra Meyer (Kaiserslautern), Alberto Marini (Milano), Einar Steingrìmsson (Goteborg), Harald Fripertinger (Graz), Henry Crapo (Inria, Rocquencourt), Pierre Leroux (Montréal), Pierre Lalonde (Montréal), Giuseppe Pirillo (Firenze), Andrea Brini (Bologna), Stéphane Ceroi (Lyon)